Part 1 and 2 of the same problem, not separate problemsRequired Information (The following information applies to the questions displayed below.] Part 1 of 2 Over a four-year perioRequired Information The following information applies to the questions displayed below.] Part 2 of 2 1.5 points Over a four-
Answer #1
1
2018 2019 2020 2021
Net sales 58000 64000 72000 88000
Cost of goods sold 30600 44600 26600 46600
Gross Profit 27400 19400 45400 41400
Gross Profit ratio
2018 47% =27400/58000
2019 30% =19400/64000
2020 63% =45400/72000
2021 47% =41400/88000
2
Cost of goods sold 2019 33720 =44600-10880
Cost of goods sold 2020 37480 =26600+10880
2018 2019 2020 2021
Net sales 58000 64000 72000 88000
Cost of goods sold 30600 33720 37480 46600
Gross Profit 27400 30280 34520 41400
Gross Profit ratio
2018 47% =27400/58000
2019 47% =30280/64000
2020 48% =34520/72000
2021 47% =41400/88000